Scampini tech is expected to generate $25 million in free cash flow next year and the FCF is expected to grow at a constant rate of 41% per year indefinitely. Scampini has no debt or preferred stock and its WACC is 10%. If scampini has $40 million shares of stock outstanding, what is the stocks value per share?
Answer #1

note : Growth rate should be 4% and not 41%.. can u please check and confirm...

using 4% below is the answer -

a Total value of the equity = free cash flow next year/(required rate - growth rate)
=25000000*104%/(10%-41%)
433333333.33
b Number of share = 40000000
c=a/b Price per share =                10.83